# LoRA-Pro: Are Low-Rank Adapters Properly Optimized?

### v2

@@ Abstract
Low-rank adaptation, also known as LoRA, has emerged as a prominent method for parameter-efficient fine-tuning of foundation models. Despite its computational efficiency, LoRA still yields inferior performance compared to full fine-tuning. In this paper, we first uncover a fundamental connection between the optimization processes of LoRA and full fine-tuning: using LoRA for optimization is mathematically equivalent to full fine-tuning using a low-rank gradient for parameter updates. And this low-rank gradient can be expressed in terms of the gradients of the two low-rank matrices in LoRA. Leveraging this insight, we introduce LoRA-Pro, a method that enhances LoRA’s performance by strategically adjusting the gradients of these low-rank matrices. This adjustment allows the low-rank gradient to more accurately approximate the full fine-tuning gradient, thereby narrowing the performance gap between LoRA and full fine-tuning. Furthermore, we theoretically derive the optimal solutions for adjusting the gradients of the low-rank matrices, applying them during fine-tuning in LoRA-Pro. We conduct extensive experiments across natural language understanding, dialogue generation, mathematical reasoning, code generation, and image classification tasks, demonstrating that LoRA-Pro substantially improves LoRA’s performance, effectively narrowing the gap with full fine-tuning. Code is publicly available at https://github.com/mrflogs/LoRA-Pro.
@@ 1 Introduction
  • We first uncover a crucial connection between LoRA and full fine-tuning in the optimization process: optimizing with LoRA is mathematically equivalent to full fine-tuning using a low-rank gradient for updating.
@@ 3 Experimental Results
Training details. To ensure a fair comparison, we align our experimental setup with that of LoRA-GA (Wang et al., 2024). By default, we fine-tune the model using the AdamW optimizer (Loshchilov & Hutter, 2019) with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay set to 0. We implement a cosine learning rate schedule with a warmup ratio of 0.03. LoRA is applied to all linear modules, excluding the embedding layer, normalization layer, and classification head. By default, we set the rank $r=8$ and $\alpha=16$.
@@ 3.1 Results on Natural Language Understanding Tasks
In this section, we evaluate our LoRA-Pro across various natural language understanding datasets. To provide a comprehensive comparison, we include several baseline methods: 1) full fine-tuning and the standard LoRA (Hu et al., 2022). 2) LoRA variants maintaining the original structure, such as rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), PiSSA (Meng et al., 2024), and LoRA-GA (Wang et al., 2024), 3) LoRA variants with modified structures, including DoRA (Liu et al., 2024) and AdaLoRA (Zhang et al., 2023).
@@
LoRA-Pro 5.86±0.06 54.23±0.79 22.76±0.35
LoRA-Pro (rank=32) 6.01±0.05 55.14±1.73 28.05±0.00
LoRA-Pro (rank=128) 5.68±0.14 56.48±0.23 34.55±2.46
@@ 3.2 Results on Large Language Models
In this section, we evaluate the performance of LoRA-Pro on large language models, focusing on dialogue generation, mathematical reasoning, and code generation capabilities. Our experimental setup follows the configuration used in LoRA-GA (Wang et al., 2024).
@@
We compare LoRA-Pro with several baselines, including full fine-tuning, LoRA (Hu et al., 2022), PiSSA (Meng et al., 2024), rsLoRA (Kalajdzievski, 2023), LoRA+(Hayou et al., 2024), DoRA(Liu et al., 2024), AdaLoRA (Zhang et al., 2023), and LoRA-GA (Wang et al., 2024). By default, we set the rank to 8 and $\alpha=16$. Following LoRA-GA (Wang et al., 2024), we initialize the scaling factor as in rsLoRA (Kalajdzievski, 2023), i.e., $s=\frac{\alpha}{\sqrt{r}}$.
Table 2 presents our experimental results, which demonstrate LoRA-Pro’s superior performance. With a rank of 8, LoRA-Pro achieves notable improvements over the original LoRA: 0.25 on MT-Bench, 11.98 on GSM8K, and 8.00 on HumanEval. When compared to the second-best PEFT method, LoRA-GA, LoRA-Pro shows consistent gains: 0.46 on GSM8K and a substantial 2.95 on HumanEval. These results validate the effectiveness of our LoRA-Pro method.
@@
To further explore the scalability of our method, we increase the rank in LoRA-Pro from 8 to 128. Our observations reveal a clear trend: as the rank increases, the performance gap between LoRA-Pro and full fine-tuning narrows rapidly. Notably, LoRA-Pro consistently outperforms LoRA-GA at the same ranks on both GSM8K and HumanEval datasets. At rank 32, LoRA-Pro surpasses LoRA-GA by 0.02 on GSM8K and 7.87 on HumanEval. This performance disparity becomes even more pronounced at rank 128, where LoRA-Pro outperforms LoRA-GA by 1.41 on GSM8K and 11.5 on HumanEval. These results demonstrate the superior scalability and effectiveness of LoRA-Pro across various rank settings.
@@ 3.3 Results on Image Classification Tasks
In this section, we assess the performance of LoRA-Pro on image classification tasks. To provide a comprehensive comparison, we compare it with several baselines: zero-shot CLIP (Radford et al., 2021), full fine-tuning, vanilla LoRA (Hu et al., 2022), rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), and DoRA (Liu et al., 2024).
@@ 3.4 Ablation Study
Table 4: Ablation study on the selection of different X in LoRA-Pro.
choice of X MT-Bench GSM8K HumanEval
Zero 5.76±0.02 53.83±1.16 22.96±1.96
Sylvester (Thm. 2.3) 5.86±0.06 54.23±0.79 22.76±0.35
Symmetry (Eq. (16)) 5.63±0.12 54.46±0.88 22.56±1.06
Ablation study on the selection of X. Based on Theorem 2.1, in LoRA-Pro, the matrix $X$ can be chosen arbitrarily. While its selection does not affect the equivalent gradient, it does influence the updates of matrices $A$ and $B$ in LoRA. Here, we conduct an ablation study on the choice of $X$.
$X=-\frac{1}{2s}B(B^{T}B)^{-1}B^{T}gA(A^{T}A)^{-1}A=-\frac{1}{2s^{2}}B(B^{T}B)^{-1}B^{T}g^{B}_{lora}(A^{T}A)^{-1}A.$ (16)
The comparison of the selection of $X$ is presented in Table 4. As shown in the table, the results obtained from different choices of $X$ are similar, and, except on MT-Bench, they outperform the other PEFT baselines listed in Table 2. However, considering the high variance in the results for both the zero solution and the symmetry solution, we ultimately select the sylvester solution as the default choice for the matrix $X$.
@@
In Figure 2, we illustrate the rank changes of matrices $A$ and $B$ from the $q$ projection of layer 9 during the training process, with rank set to 8 and 32, resipectively. We observed that, although $A$ and $B$ do not initially satisfy the full rank assumption (with matrix $B$ initialized as a zero matrix), both matrices achieve full rank after the first update step. The rank behavior of $A$ and $B$ in other layers also exhibits similar results.
@@
Table 5: We compare LoRA, LoRA-Pro, and Full Fine-Tuning in terms of memory cost, training time, and performance on the MT-Bench, GSM8K, and HumanEval datasets. Memory cost is measured using a single A6000 GPU with a batch size of 1. Training time is recorded on the WizardLM dataset using 8 A100 GPUs with DeepSpeed ZeRO-2 stage optimization.
Memory Cost Training Time MT-Bench GSM8K HumanEval
Full FT $>$ 48 GB 2h 33min 5.30±0.11 59.36±0.85 35.31±2.13
LoRA 22.26 GB 1h 22min 5.61±0.10 42.08±0.04 14.76±0.17
LoRA-Pro 23.05 GB 1h 23min 5.86±0.06 54.23±0.79 22.76±0.35
Memory footprint and training time. Here, we evaluate the additional costs associated with using LoRA-Pro compared to LoRA. We primarily focus on comparing the differences in memory cost and training time between LoRA-Pro, LoRA, and full fine-tuning. The results of the experiments are presented in Table 5. We measure memory cost using a single A6000 GPU with a batch size of 1. Training time is recorded on the WizardLM dataset using 8 A100 GPUs with DeepSpeed (Rasley et al., 2020) ZeRO-2 stage optimization.
The results are shown in Table 5. From the table, we observe the following: 1) LoRA-Pro requires approximately 0.8GB more GPU memory compared to LoRA. This difference likely stems from the need to compute $B^{T}B$, $AA^{T}$, and their inverses during the calculation of the optimal gradients. 2) Surprisingly, the training time for LoRA-Pro is nearly identical to that of LoRA, with only about a 1% increase in additional training time. We attribute this to the fact that the matrices $A$ and $B$ in LoRA are low-rank. Consequently, the extra computations required by LoRA-Pro (such as matrix inversion and the calculation of $X$) are performed on small $r\times r$ matrices, resulting in negligible computational overhead.
@@ 4 Related Work
Parameter-Efficient Fine-Tuning. Given the huge size of foundation models, recent research has focused on developing parameter-efficient fine-tuning methods (Hu et al., 2022; Liu et al., 2024; Ding et al., 2023; Houlsby et al., 2019; Liu et al., 2023; Lester et al., 2021). These methods aim to reduce the cost of fine-tuning by adjusting only a small portion of the model’s parameters. Generally, these methods fall into two main categories. The first category is adapter tuning (Houlsby et al., 2019; Sung et al., 2022; He et al., 2021; Zhang et al., 2024; Bapna & Firat, 2019; Hu et al., 2022), which involves inserting small neural network modules, called adapters, into specific layers of the model. During fine-tuning, we keep the model frozen and only fine-tune the lightweight adapter modules, significantly reducing the memory footprint for fine-tuning. The second category is prompt tuning (Lester et al., 2021; Zhou et al., 2022; Li & Liang, 2021; Liu et al., 2022). Prompt tuning adapts the models to specific tasks by adding specially designed prompts or learnable tokens to the input data, rather than directly modifying the internal parameters of foundation models. In this paper, we focus on LoRA (Hu et al., 2022), a prominent method within the realm of adapter tuning.
Low-Rank Adaptation. Low-rank adaptation, initially referred to as LoRA (Hu et al., 2022), has evolved into a broad category encompassing parameter-efficient fine-tuning methods based on low-rank approximations (Hu et al., 2022; Liu et al., 2024; Hayou et al., 2024; Kalajdzievski, 2023; Zhang et al., 2023; Kopiczko et al., 2024; Hyeon-Woo et al., 2022; Zhang & Pilanci, 2024; Wang et al., 2024; Zhao et al., 2024; Wang et al., 2024). LoRA (Hu et al., 2022) assumes that the changes in the weights of pre-trained models exhibit a low-rank structure. Consequently, it re-parameterizes these changes as the product of low-rank matrices, thereby reducing the cost during fine-tuning.
@@ Appendix C Optimization Algorithms
In this section, we present the pseudo-codes for implementing our LoRA-Pro method using the SGD (Sutskever et al., 2013) and AdamW (Loshchilov & Hutter, 2019) optimizers. The details are provided in Algorithm 1, Algorithm 2, and Algorithm 3, respectively.
@@
LoRA-Pro with AdamW optimizer. In AdamW optimizer, the implementation becomes more complex. We offer two scalable implementation modes for LoRA-Pro with AdamW: efficient mode and full mode. We use the efficient mode by default. In efficient mode, the memory usage of LoRA-Pro’s optimizer remains the same as that of standard LoRA, but this mode tends to be less stable. In full mode, LoRA-Pro aligns more closely with full fine-tuning during the optimization process, resulting in higher memory usage but improved stability.
In efficient mode, we only need to modify the gradients of matrices $A$ and $B$. Then, we separately compute the first- and second-order gradient information for $A$ and $B$ before performing the update. The detailed algorithm is shown in Algorithm 2. In this mode, memory usage remains consistent with that of LoRA (Hu et al., 2022), delivering better results, though not as stable as full mode.
In full mode, we aim to closely approximate full fine-tuning during optimization. Several modifications are necessary. Firstly, in order to mimic full fine-tuning, after adjusting the gradients of matrices $A$ and $B$, we need to compute the equivalent gradient,
@@
Ablation study on the AdamW optimizer. We conducted experiments to compare the memory cost differences between the two implementations of the AdamW optimizer, as well as their performance on MT-Bench, GSM8K, and HumanEval. The results of the study are presented in Table 7. As shown in the table, the full mode achieves better performance compared to the efficient mode. We believe this is because the full mode more closely aligns with full fine-tuning during the optimization process. However, as a trade-off, full mode incurs a higher GPU memory cost due to the calculation of the first- and second-order momentum information for the equivalent gradients $\tilde{g}$. Despite this, its memory usage is still significantly lower than that of full fine-tuning (which theoretically requires 56GB of memory). Therefore, we use efficient mode as the default setting for LoRA-Pro.
Table 7: Comparison of the AdamW optimizer implementations with LoRA-Pro. We present two types of AdamW optimizers for LoRA-Pro: the efficient mode, which maintains the same memory cost as the vanilla LoRA optimizer, and the full mode, which is more aligned with full fine-tuning and offers greater stability during training.
Memory Cost MT-Bench GSM8K HumanEval
Full FT $>$ 48 GB 5.30±0.11 59.36±0.85 35.31±2.13
LoRA-Pro (Efficient) 23.05 GB 5.86±0.06 54.23±0.79 22.76±0.35
LoRA-Pro (Full) 43.87 GB 6.05±0.25 54.06±0.46 23.98±0.35
Algorithm 2 LoRA-Pro with AdamW optimizer (Efficient Mode)
0:  Given initial learning rate $\gamma$, scaling factor $s$, and $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8},\lambda\in\mathbb{R}$
1:  Initialize time step $t\leftarrow 0$, low-rank matrices $A_{0}\in\mathbb{R}^{r\times n}$ and $B_{0}\in\mathbb{R}^{m\times r}$, first momentum $m^{A}_{0}\in\mathbb{R}^{r\times n}$, $m^{B}_{0}\in\mathbb{R}^{m\times r}$, second momentum $v^{A}_{t}\in\mathbb{R}^{r\times n}$, $v^{B}_{t}\in\mathbb{R}^{m\times r}$
2:  repeat
3:     $t\leftarrow t+1$
[…]
8:     $g^{B}=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX$
9:     $m^{A}_{t}\leftarrow\beta_{1}m^{A}_{t-1}+(1-\beta_{1})g^{A}$
10:     $v^{A}_{t}\leftarrow\beta_{2}v^{A}_{t-1}+(1-\beta_{2})g^{A}*g^{A}$
11:     $\hat{m}^{A}_{t}\leftarrow\frac{m_{t}^{A}}{1-\beta_{1}^{t}}$
12:     $\hat{v}^{A}_{t}\leftarrow\frac{v_{t}^{A}}{1-\beta_{2}^{t}}$
13:     $\tilde{g}^{A}\leftarrow\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
14:     $A_{t}\leftarrow(1-\lambda\gamma)A_{t-1}$
15:     $A_{t}\leftarrow A_{t}-\gamma\tilde{g}^{A}$
16:     $m^{B}_{t}\leftarrow\beta_{1}m^{B}_{t-1}+(1-\beta_{1})g^{B}$
17:     $v^{B}_{t}\leftarrow\beta_{2}v^{B}_{t-1}+(1-\beta_{2})g^{B}*g^{B}$
18:     $\hat{m}^{B}_{t}\leftarrow\frac{m_{t}^{B}}{1-\beta_{1}^{t}}$
19:     $\hat{v}^{B}_{t}\leftarrow\frac{v_{t}^{B}}{1-\beta_{2}^{t}}$
20:     $\tilde{g}^{B}\leftarrow\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
21:     $B_{t}\leftarrow B_{t-1}-\gamma\tilde{g}^{B}$
22:  until stopping criterion is met
23:  return  optimized parameters $A_{t}$ and $B_{t}$
Algorithm 3 LoRA-Pro with AdamW optimizer (Full Mode)

### v3

@@ Abstract
Low-rank adaptation, also known as LoRA, has emerged as a prominent method for parameter-efficient fine-tuning of foundation models. Despite its computational efficiency, LoRA still yields inferior performance compared to full fine-tuning. In this paper, we first uncover a fundamental connection between the optimization processes of LoRA and full fine-tuning: using LoRA for optimization is mathematically equivalent to full fine-tuning using a low-rank gradient for parameter updates. And this low-rank gradient can be expressed in terms of the gradients of the two low-rank matrices in LoRA. Leveraging this insight, we introduce LoRA-Pro, a method that enhances LoRA’s performance by strategically adjusting the gradients of these low-rank matrices. This adjustment allows the low-rank gradient to more accurately approximate the full fine-tuning gradient, thereby narrowing the performance gap between LoRA and full fine-tuning. Furthermore, we theoretically derive the optimal solutions for adjusting the gradients of the low-rank matrices, applying them during fine-tuning in LoRA-Pro. We conduct extensive experiments across natural language understanding, dialogue generation, mathematical reasoning, code generation, and image classification tasks, demonstrating that LoRA-Pro substantially improves LoRA’s performance, effectively narrowing the gap with full fine-tuning. Our code is publicly available at https://github.com/mrflogs/LoRA-Pro.
@@ 1 Introduction
  • We first uncover a crucial connection between LoRA and full fine-tuning in optimization process: optimizing with LoRA is mathematically equivalent to full fine-tuning using a low-rank gradient for updating.
@@ 3 Experimental Results
Training details. To ensure a fair comparison, we align our experimental setup with that of LoRA-GA (Wang et al., 2024a). By default, we fine-tune the model using the AdamW optimizer (Loshchilov & Hutter, 2019) with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay set to 0. We implement a cosine learning rate schedule with a warmup ratio of 0.03. LoRA is applied to all linear modules, excluding the embedding layer, normalization layer, and classification head. By default, we set the rank $r=8$ and $\alpha=16$.
@@ 3.1 Results on Natural Language Understanding Tasks
In this section, we evaluate our LoRA-Pro across various natural language understanding datasets. To provide a comprehensive comparison, we include several baseline methods: 1) full fine-tuning and the standard LoRA (Hu et al., 2022). 2) LoRA variants maintaining the original structure, such as rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), PiSSA (Meng et al., 2024), and LoRA-GA (Wang et al., 2024a), 3) LoRA variants with modified structures, including DoRA (Liu et al., 2024) and AdaLoRA (Zhang et al., 2023).
@@
LoRA-Pro 5.72±0.03 57.57±0.50 22.97±0.35
LoRA-Pro (rank=32) 5.57±0.51 57.97±0.50 26.63±0.35
LoRA-Pro (rank=128) 5.67±0.11 61.08±0.19 30.28±0.93
@@ 3.2 Results on Large Language Models
In this section, we evaluate the performance of LoRA-Pro on large language models, focusing on dialogue generation, mathematical reasoning, and code generation capabilities. Our experimental setup follows the configuration used in LoRA-GA (Wang et al., 2024a).
@@
We compare LoRA-Pro with several baselines, including full fine-tuning, LoRA (Hu et al., 2022), PiSSA (Meng et al., 2024), rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023), and LoRA-GA (Wang et al., 2024a). By default, we set the rank to 8 and $\alpha=16$. Following LoRA-GA (Wang et al., 2024a), we initialize the scaling factor as in rsLoRA (Kalajdzievski, 2023), i.e., $s=\frac{\alpha}{\sqrt{r}}$.
Table 2 presents our experimental results, which demonstrate LoRA-Pro’s superior performance. With a rank of 8, LoRA-Pro achieves notable improvements over the original LoRA: 0.11 on MT-Bench, 15.49 on GSM8K, and 8.21 on HumanEval. When compared to the second-best PEFT method, LoRA-GA, LoRA-Pro shows consistent gains: 3.97 on GSM8K and a substantial 3.16 on HumanEval. These results validate the effectiveness of our LoRA-Pro method.
@@
To further explore the scalability of our method, we increase the rank in LoRA-Pro from 8 to 128. Our observations reveal a clear trend: as the rank increases, the performance gap between LoRA-Pro and full fine-tuning narrows rapidly. Notably, LoRA-Pro consistently outperforms LoRA-GA at the same ranks on both GSM8K and HumanEval datasets. At rank 32, LoRA-Pro surpasses LoRA-GA by 2.85 on GSM8K and 6.45 on HumanEval. This performance disparity becomes even more pronounced at rank 128, where LoRA-Pro outperforms LoRA-GA by 6.01 on GSM8K and 7.23 on HumanEval. These results demonstrate the superior scalability and effectiveness of LoRA-Pro across various rank settings.
@@ 3.3 Results on Image Classification Tasks
In this section, we assess the performance of LoRA-Pro on image classification tasks. To provide a comprehensive comparison, we compare it with several baselines: zero-shot CLIP (Radford et al., 2021), full fine-tuning, vanilla LoRA (Hu et al., 2022), rsLoRA (Kalajdzievski, 2023), LoRA+ (Hayou et al., 2024), DoRA (Liu et al., 2024), and LoRA-GA (Wang et al., 2024a).
@@ 3.4 Ablation Study
In Figure 2, we illustrate the rank changes of matrices $A$ and $B$ from the $q$ projection of layer 9 during the training process, with rank set to 8 and 32, respectively. We observed that, although $A$ and $B$ do not initially satisfy the full rank assumption (with matrix $B$ initialized as a zero matrix), both matrices achieve full rank after the first update step. The rank behavior of $A$ and $B$ in other layers also exhibits similar results.
@@
Table 4: We compare LoRA, LoRA-Pro, and Full Fine-Tuning in terms of memory cost, training time, and performance on the MT-Bench, GSM8K, and HumanEval datasets. Memory cost is measured using a single A6000 GPU with a batch size of 1. Training time is recorded on the MetaMathQA dataset using 8 A100 GPUs with DeepSpeed ZeRO-2 stage optimization.
Method Memory Cost Training Time MT-Bench GSM8K HumanEval
Full FT $\sim 8\times 40$ GB 4h 20min 5.30±0.11 59.36±0.85 35.31±2.13
LoRA $\sim 8\times 17$ GB 1h 30min 5.61±0.10 42.08±0.04 14.76±0.17
LoRA-GA $\sim 8\times 17$ GB 1h 31min 5.95±0.16 53.60±0.30 19.81±1.46
LoRA-Pro $\sim 8\times 21$ GB 1h 41min 5.72±0.03 57.57±0.50 22.97±0.35
Memory footprint and training time. Here, we evaluate the additional costs associated with using LoRA-Pro compared to LoRA. We primarily focus on comparing the differences in memory cost and training time between LoRA-Pro, LoRA, and full fine-tuning. The results of the experiments are presented in Table 4. We measure memory cost using 8 A6000 GPUs with a batch size of 1. Training time is recorded on the WizardLM dataset using 8 A100 GPUs with DeepSpeed (Rasley et al., 2020) ZeRO-2 stage optimization.
The results are shown in Table 4. From the table, we observe the following: 1) LoRA-Pro requires approximately 4GB more GPU memory compared to LoRA. This difference likely stems from the need to compute $B^{T}B$, $AA^{T}$, and their inverses during the calculation of the optimal gradients. 2) Surprisingly, the training time for LoRA-Pro is nearly identical to that of LoRA, with only about 10 minutes increase in additional training time. We attribute this to the fact that the matrices $A$ and $B$ in LoRA are low-rank. Consequently, the extra computations required by LoRA-Pro (such as matrix inversion and the calculation of $X$) are performed on small $r\times r$ matrices, making the extra computational overhead manageable.
@@
Figure 3: Training loss curves of LoRA, LoRA-GA, LoRA-Pro, and Full Fine-tuning on WizardLM, MetaMathQA, and CodeFeedback.
Training curves of LoRA-Pro. In this part, we present the training loss curves for LoRA, LoRA-GA, LoRA-Pro, and Full Fine-tuning across WizardLM, MetaMathQA, and CodeFeedback datasets. As illustrated in Figure 3, LoRA-Pro demonstrates a faster convergence speed compared to LoRA and LoRA-GA. Furthermore, LoRA-Pro achieves a lower final loss value upon convergence, indicating its improved efficiency and effectiveness.
@@ 4 Related Work
Parameter-Efficient Fine-Tuning. Given the huge size of foundation models, recent research has focused on developing parameter-efficient fine-tuning methods (Hu et al., 2022; Liu et al., 2024; Ding et al., 2023; Houlsby et al., 2019; Liu et al., 2023; Lester et al., 2021; Wang et al., 2024c). These methods aim to reduce the cost of fine-tuning by adjusting only a small portion of the model’s parameters. Generally, these methods fall into two main categories. The first category is adapter tuning (Houlsby et al., 2019; Sung et al., 2022; He et al., 2021; Zhang et al., 2024; Bapna & Firat, 2019; Hu et al., 2022), which involves inserting small neural network modules, called adapters, into specific layers of the model. During fine-tuning, we keep the model frozen and only fine-tune the lightweight adapter modules, significantly reducing the memory footprint for fine-tuning. The second category is prompt tuning (Lester et al., 2021; Zhou et al., 2022; Li & Liang, 2021; Liu et al., 2022; Wang et al., 2023; 2024b; Liang et al., 2024). Prompt tuning adapts the models to specific tasks by adding specially designed prompts or learnable tokens to the input data, rather than directly modifying the internal parameters of foundation models. In this paper, we focus on LoRA (Hu et al., 2022), a prominent method within the realm of adapter tuning.
Low-Rank Adaptation. Low-rank adaptation, initially referred to as LoRA (Hu et al., 2022), has evolved into a broad category encompassing parameter-efficient fine-tuning methods based on low-rank approximations (Hu et al., 2022; Liu et al., 2024; Hayou et al., 2024; Kalajdzievski, 2023; Zhang et al., 2023; Kopiczko et al., 2024; Hyeon-Woo et al., 2022; Zhang & Pilanci, 2024; Wang et al., 2024a; Zhao et al., 2024; Wang et al., 2024a). LoRA (Hu et al., 2022) assumes that the changes in the weights of pre-trained models exhibit a low-rank structure. Consequently, it re-parameterizes these changes as the product of low-rank matrices, thereby reducing the cost during fine-tuning.
@@ 5 Conclusion
Limitations. LoRA-Pro still have some limitations: (1) LoRA-Pro still adheres to LoRA’s assumption that $\Delta W$ is of low rank. However, this assumption may break down in cases of pre-training or when there is a large amount of fine-tuning data, potentially leading to suboptimal results. (2) So far, we have only applied LoRA-Pro to variants that have a structure similar to LoRA. It currently cannot be applied to structurally different LoRA models, such as DoRA (Liu et al., 2024) or FLoRA (Wen & Chaudhuri, 2024). We plan to explore these directions in future research.
Acknowledgement
This work was funded by the National Natural Science Foundation of China under Grants (62276256, U2441251) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001). We thank Jie Cheng and Yongcan Yu for providing computational resources critical to this work.
@@ Appendix C Optimization Algorithms
In this section, we present the pseudo-codes for implementing our LoRA-Pro method using the SGD (Sutskever et al., 2013) and AdamW (Loshchilov & Hutter, 2019) optimizers. The details are provided in Algorithm 1 and Algorithm 2, respectively.
@@
LoRA-Pro with AdamW optimizer. In AdamW optimizer, the implementation becomes more complex. We aim to closely approximate full fine-tuning during optimization. Several modifications are necessary. Firstly, in order to mimic full fine-tuning, after adjusting the gradients of matrices $A$ and $B$, we need to compute the equivalent gradient,
@@
Algorithm 2 LoRA-Pro with AdamW optimizer
@@
Appendix D Additional Experiments
D.1 Ablation Study of the Selection of X
Based on Theorem 2.1, in LoRA-Pro, the matrix $X$ can be chosen arbitrarily. While its selection does not affect the equivalent gradient, it does influence the updates of matrices $A$ and $B$ in LoRA. Here, we conduct an ablation study on the choice of $X$.
$X=-\frac{1}{2s}(B^{T}B)^{-1}B^{T}gA(A^{T}A)^{-1}=-\frac{1}{2s^{2}}(B^{T}B)^{-1}B^{T}g^{B}_{lora}(A^{T}A)^{-1}.$ (58)
The comparison of the selection of $X$ is presented in Table 6. As shown in the table, the choice of $X$ significantly impacts LoRA-Pro’s performance, particularly evident in the GSM8K dataset. Different $X$ selections influence the subspaces of $A$ and $B$, ultimately affecting the optimization process described in Theorem 2.1. Our experiments demonstrate that the Sylvester solution consistently outperforms both the zero and symmetry solutions across all three evaluation tasks. We attribute the superior performance of the Sylvester solution to its ability to select subspaces for $A$ and $B$ that enable faster gradient descent (i.e., maximizing the approximation between the modified gradients $g^{A},g^{B}$ and the LoRA gradients $g^{A}_{lora},g^{B}_{lora}$).
Table 6: Ablation study on the selection of different X in LoRA-Pro.
Choice of X MT-Bench GSM8K HumanEval
Zero 5.58±0.14 31.74±0.69 17.28±0.35
Symmetry (Eq. (58)) 5.71±0.11 42.81±0.62 17.88±0.35
Sylvester (Thm. 2.3) 5.72±0.03 57.57±0.50 22.97±0.35
D.2 Visualization of Differences Between Equivalent Gradients and Full Gradients
In this section, we fine-tune Llama-2-7B on the MetaMathQA100k dataset and visualize the discrepancies between the equivalent gradients of LoRA and LoRA-Pro and the full gradients during training, i.e., the differences before and after gradient adjustments. We present visualizations for different optimization modules, including the Q, K, V, O, Up, Down, and Gate layers, and provide results for these modules across the shallow (1), medium (15), and deep (31) layers of Llama-2-7B.
The results are shown in Figure 4. From the figure, we can draw the following conclusions:
  • After gradient adjustments in LoRA-Pro, we observe a significant reduction in the distance between the equivalent gradients and the full gradients.
  • In certain layers, the discrepancy between LoRA’s equivalent gradients and the full gradients continues to increase (e.g., Layer 1 O, Up, Gate projections; Layer 15 Up and Gate projections; and Layer 31 O projection). However, in these layers, the discrepancy for LoRA-Pro remains stable, indicating that LoRA-Pro can consistently align with the full gradients during training, preventing the model from settling into sub-optimal solutions.
  • In deep layers, the discrepancy between equivalent gradients and full gradients decreases as training progresses, whereas in shallow and medium layers, the discrepancy first increases and then stabilizes. The cause of this phenomenon is not yet clear, and we plan to investigate it further in future research.
These findings highlight that LoRA-Pro effectively reduces the distance between LoRA and full gradients during training and ensures continuous alignment with full gradients, underscoring the efficacy of LoRA-Pro.
Figure 4: Visualization of the differences between the equivalent gradients of LoRA, LoRA-Pro, and the full-parameter gradients during training, i.e., $\|\tilde{g}-g\|_{F}$. The rows illustrate the differences across various modules, including Q, K, V, O, Up, Down, and Gate. The columns show the differences at different depths, categorized as shallow (1), medium (15), and deep layers (31).
D.3 Experiments Results with Different Learning Rates
To demonstrate the effectiveness of LoRA-Pro, we evaluated its performance on GSM8K under learning rates of 1e-5 and 5e-5, comparing it with LoRA and LoRA-GA. The results, presented in Table 7, show that LoRA-Pro maintains its advantages under both learning rates, highlighting its robustness to variations in learning rate.
Table 7: Performance comparison of LoRA, LoRA-GA, LoRA-Pro on GSM8K with learning rates 1e-5, 2e-5, and 5e-5.
GSM8K LoRA LoRA-GA LoRA-Pro
1e-5 36.65±0.82 50.25±0.62 56.48±0.27
2e-5 42.08±0.04 53.60±0.30 57.57±0.50
5e-5 46.41±0.16 52.89±0.19 58.76±1.86
D.4 Additional Experiments on Latest Models
Table 8: Performance comparison of LoRA, LoRA-GA, and LoRA-Pro with Llama-2-7B and Llama-3.1-8B.
GSM8K LoRA LoRA-GA LoRA-Pro
Llama-2-7B 42.08±0.04 53.60±0.30 57.57±0.50
Llama-3.1-8B 71.04±0.26 72.20±1.15 75.49±0.42
To further demonstrate the effectiveness of LoRA-Pro, we conducted additional experiments using the latest model, LLaMA-3.1-8B (Dubey et al., 2024). We fine-tuned the model using these three methods,LoRA, LoRA-GA, and LoRA-Pro,on the MetaMath100k dataset and evaluated its performance on the GSM8k dataset. All results are averaged over three different random seeds.
As shown in Table 8, LoRA-Pro demonstrates a clear advantage over both LoRA and LoRA-GA when applied to the LLaMA-3.1-8B model, further highlighting its effectiveness.